OpenStudy (anonymous):

what is the slope intercept form of the following linear equation y-3=2(x-1/2)

jimthompson5910 (jim_thompson5910):

y-3=2(x-1/2) y-3=2(x)+2(-1/2) y-3=2x-1 y=2x-1+3 y=2x-2

OpenStudy (anonymous):

I have y=2x+2 as an answer but not 2x-2

jimthompson5910 (jim_thompson5910):

oops should be y = 2x+2

jimthompson5910 (jim_thompson5910):

yeah made a typo, you have it right